# From Infants to AI: Incorporating Infant-like Learning in Models Boosts Efficiency and Generalization in Learning Social Prediction Tasks

### v1

@@ Introduction
The learning of early concepts comes together with an initial understanding of aspects of the concepts’ meaning, in terms of their implications, causality, or use for predicting likely future events. For example, in learning to identify hands, infants also learn that hands cause objects to move and change location (?, ?, ?). All this learning is obtained in many cases with little or no supervision, and from relatively few examples, compared with current models (?, ?). In learning about objects and human-object interactions, early learned concepts are often used in the process of acquiring additional related concepts. For example, the relation of in front/behind appears to be a prerequisite to the subsequent learning of containment (?, ?). Learning to identify the direction of gaze is used to predict future actions (?, ?), and it later plays a role in the development of communication and language (?, ?).
@@
BERT is a language model trained on a substantial amount of human language. However, we ensured that this prior knowledge did not affect our results by confirming that the initial performance on our tasks was not significantly different from the chance level.
@@ Results
Two task types were generated: Task 1 (T1): A simpler task where the test data use the same actors as in the training data but introduce new target objects. Task 2 (T2): A more challenging task where the test data include entirely new actors and new target objects.
The networks were first trained on T1 and then further fine-tuned under one of two conditions: T1-T1: The same simple task (T1) with additional data. This condition served as a control to assess the effect of additional data without introducing new challenges. T1-T2: The more challenging generalization task (T2). Each task used 640 training examples and 160 test examples. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
Figure 4: Input Data for Experiments 2 and 3: The five rightmost frames were used in Experiment 2, while the full seven frames are used in Experiment 3.
@@
The results for both models are shown in Figure 5, comparing the performance under the following conditions: 1. Learning the initial task (T1). 2. Further training on the same task with additional data (T1-T1). 3. Generalization to a new task (T1-T2).
The cognitive network outperformed the naive network in both accuracy and learning speed across all tasks: T1-T1 Condition: The cognitive network achieved near-perfect accuracy in the first run and 100% accuracy almost immediately in the second run. The naive network stabilized at 58% accuracy after 1000 epochs in the first run and reached 75% accuracy after 200 epochs in the second run. T1-T2 Condition: For the challenging generalization task, the cognitive network achieved perfect accuracy after just a few epochs. The naive network, however, reached only  65% accuracy, performing worse than in the T1-T1 condition.
@@
Figure 5: Experiment 2: Generalization Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1), further training on Task 1 (T1-T1), and further training on Task 2 (T1-T2), with SEM included.
@@
We address the following question: How is the additional information provided to the cognitive model learned? In earlier experiments, the actor’s type (animate/inanimate) was explicitly provided to the cognitive network. Here, both networks were tasked with autonomously learning the actor’s type from the data. Inspired by evidence that self-propelled motion is a strong indicator of animacy in infants (?, ?), the experiment introduced additional frames to enable the learning of this concept.
@@
The naive network was tested on the 7-frame paradigm with progressively larger datasets, starting from 1280 training and 320 test examples, doubling in size at each step. The cognitive network, in contrast, learned actor types using smaller datasets of 320 training and 80 test examples, which proved sufficient. For the 5-frame task, a dataset of 640 training and 160 test examples was used. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
@@
The cognitive network learned the first task of classifying the actor’s type within 20 epochs, achieving perfect accuracy (Figure 6). For the subsequent task, the cognitive network performs the 5-frame paradigm discussed earlier, so the results for this stage are identical to those of the 5-frame test. The performance of the naive network is shown in Figure 7, alongside the cognitive network’s results from the 5-frame stage. For the smallest dataset, the naive network achieved only chance-level performance. When the dataset size was doubled, its accuracy increased to approximately 60%. With the largest dataset, accuracy improved to 88%. In comparison, the cognitive network achieved near-perfect accuracy on the ensuing 5-frame test, even with the smallest dataset.
@@ Discussion
The current study examined a basic characteristic of early human learning, by modeling some aspects of how early-acquired concepts are used in the learning of subsequent concepts and comparing the results with standard deep network modeling. The results show that the use of early concepts in the learning of new concepts can lead to higher accuracy and more efficient learning. The results of comparing generalization to new tasks also suggest that the integration of early and new concepts during the learning process can shape the representation of the concepts acquired by the model.
@@
Current large models are trained on huge data sets, and they strive to create so-called foundational models, which can deal with a broad range of tasks. Such models have achieved impressive results. However, multiple studies have shown examples where the learning by current large models appears to be different from human leaning; they sometimes ‘show failures on surprisingly trivial problems’ (?, ?), or ‘show basic errors in understanding that would not be expected even in non-expert humans’ (?, ?). It is possible that shortcomings of this kind will be overcome by additional data and training. Alternatively, it may prove useful to combine current large models and human learning in some useful manner. For example, it may be possible to learn early concepts by adapting methods similar to self-supervised learning of early concepts in humans, and use early-learned concepts in the learning of more complex ones. If such training proves feasible, it would be of interest to test the effects of making the model more human-like both on its performance and on its similarity to human behavior in different visual tasks.

### v2

@@ Introduction
The learning of early concepts comes together with an initial understanding of aspects of the concepts’ meaning, in terms of their implications, causality, or use for predicting likely future events. For example, in learning to identify hands, infants also learn that hands cause other objects to move and change location (?, ?, ?). All this learning is obtained in many cases with little or no supervision, and from relatively few examples, compared with current models (?, ?). In learning about objects and human-object interactions, early learned concepts are often used in the process of acquiring additional related concepts. For example, the relation of in front/behind appears to be a prerequisite to the subsequent learning of containment (?, ?). Learning to identify the direction of gaze is used to predict future actions (?, ?), and it later plays a role in the development of communication and language (?, ?).
@@
We used large models (BLIP, BERT) in our cognitive model to reflect the complexity of the infant brain—infants, too, are born with a sophisticated neural architecture. We fine-tuned all layers in both models for our specific tasks. To ensure that the models’ pre-trained linguistic knowledge did not drive the observed effects, we also ran a version of the experiment using binary representations (e.g., 0/1 vectors) instead of natural language and obtained similar results. Furthermore, we confirmed that their prior knowledge did not bias the results: the models’ initial performance on our tasks was not significantly different from chance level.
@@ Results
Two task types were generated: Task 1 (T1): A simpler task where the test data use the same actors as in the training data but introduce new target objects. Task 2 (T2): A more challenging task because the test data include entirely new actors and new target objects.
The networks were first trained on T1 and then further fine-tuned under one of two conditions: T1-T1: The same simple task (T1) with additional data. This condition served as a control to assess the effect of additional data without introducing new challenges. T1-T2: The more challenging generalization task (T2). Each task used 640 training examples and 160 test examples. Both networks were trained for 1000 epochs. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged.
Figure 4: Input Data for Experiments 2 and 3: The five rightmost frames are used in Experiment 2, while the full seven frames are used in Experiment 3.
@@
The results for both models are shown in Figure 5, comparing the performance under the following conditions: 1. Learning the initial task (T1). 2. Retraining on the same task with additional data (T1-T1). 3. Generalization to a new task (T1-T2). Only the first 300 epochs are shown, as there was no significant change afterward. The cognitive network outperformed the naive network in both accuracy and learning speed across all tasks: T1-T1 Condition: The cognitive network achieved near-perfect accuracy in the first run and 100% accuracy almost immediately in the second run. The naive network stabilized at 58% accuracy after 1000 epochs in the first run and reached 75% accuracy after 200 epochs in the second run. T1-T2 Condition: For the challenging generalization task, the cognitive network achieved perfect accuracy after just a few epochs. The naive network, however, reached only  65% accuracy, performing worse than in the T1-T1 condition.
@@
Figure 5: Experiment 2: Generalization Results. Average test accuracy of the naive and cognitive networks on Task 1 (T1), retraining on Task 1 (T1-T1), and retraining on Task 2 (T1-T2), with SEM included.
@@
We address the following question: How is the additional information provided to the cognitive model learned? In earlier experiments, the actor’s type (animate/inanimate) was explicitly provided to the cognitive network. In the current experiment, however, the same information is available to both models. Inspired by evidence that self-propelled motion is a strong indicator of animacy in infants (?, ?), the experiment introduced additional frames to enable the learning of this concept.
@@
The naive network was tested on the 7-frame paradigm with progressively larger datasets, starting from 640 training and 160 test examples, doubling in size at each step. The cognitive network, in contrast, learned actor types using smaller datasets of 320 training and 80 test examples, which proved sufficient. For the 5-frame task, a dataset of 640 training and 160 test examples was used. Each experiment was repeated 15 times with randomly generated sequences, and the results were averaged. Both networks were trained for 300 epochs due to the larger datasets.
@@
The cognitive network learned the first task of classifying the actor’s type within 20 epochs, achieving perfect accuracy (Figure 6). For the subsequent task, the cognitive network performs the 5-frame paradigm discussed earlier, so the results for this stage are identical to those of the 5-frame test. The performance of the naive network is shown in Figure 7, alongside the cognitive network’s results from the 5-frame stage. For the smallest dataset, the naive network achieved approximately 60% accuracy. When the dataset size was doubled, its accuracy increased to 74%. With the largest dataset, accuracy improved to 85%. In comparison, the cognitive network achieved near-perfect accuracy on the same 5-frame test, even with the smallest dataset.
@@ Discussion
The current study examined a basic characteristic of early human learning, by modeling some aspects of how early-acquired concepts are used in the learning of subsequent concepts and comparing the results with standard deep network modeling. The results show that the use of early concepts in the learning of new concepts can lead to higher accuracy and more efficient learning. The results of comparing generalization to new tasks also suggest that the integration of early and new concepts during the learning process can shape the representation of the concepts acquired by the model. Although the dataset is simplified compared to natural images, prior work has shown that infants succeed in tasks with similarly simplified stimuli (?, ?).
@@
Current large models are trained on huge data sets, and they strive to create so-called foundational models, which can deal with a broad range of tasks. Such models have achieved impressive results. However, multiple studies have shown examples where the learning by current large models appears to be different from human leaning; they sometimes ‘show failures on surprisingly trivial problems’ (?, ?), or ‘show basic errors in understanding that would not be expected even in non-expert humans’ (?, ?). It is possible that shortcomings of this kind will be overcome by additional data and training. Alternatively, it may prove useful to combine current large models and human learning in some useful manner. For example, it may be possible to learn early concepts by adapting methods similar to self-supervised learning of early concepts in humans, and use early-learned concepts in the learning of more complex ones. If such training proves feasible, it would be of interest to test the effects of making the model more human-like both on its performance, as well as its similarity to human behavior in different visual tasks.
Acknowledgments
ST acknowledges support from the Ariane de Rothschild Women Doctoral Program. SU acknowledges the Weizmann Institute of Science for its support.